If among the 5 people, there are two women and three men. What is the probability that a 5 person committee will be made up of a woman and two men?
Answer #1

There will some correction in the question:a woman and two men, it means it needs to form 3 person committee instead of 5.

Solution

Number of ways to select 3 persons out of 5=5C3=10

now

there are two women and three men, out of that we need to select 3 person committee made up of a woman and two men.

So,

Number of ways to select a woman and two men=2C1*3C2 =2*3=6

Required Probability=6/10=0.6

Ans: 0.6